OpenStudy (bookie2cute):

what is 2n+(7n+8)= n+8 what the answer is ??

OpenStudy (anonymous):

We need to combine like terms. So first off we add 7n + 2n...what is 7+2?

OpenStudy (anonymous):

You are offline so I will just walk you through the entire process. 2n + 7n = 9n 9n + 8 = n + 8 Subtract n from both sides 8n + 8 = 8 subtract 8 8n = 0 Divide by 8 n = 0

OpenStudy (bookie2cute):

awwww thank u so much